The value of a car is inversely proportional to its age If a car is worth $4900 when it is 4 years old, how old will it be when is worth $2,800?
Answer by Cromlix(4381)

The value of a car is inversely
proportional to its age
V = k/A
First find value of k
4900 = k/4
k = 4900 x 4
k = 19600
Formula:
V = 19600/A
2800 = 19600/A
A = 19600/2800
A = 7 years
Hope this helps:-)